rxcklesskaisher:

Dawn is making an abstract painting with two triangles. The dimensions of the painting are shown below: Two triangles share the same height of 8 inches. One triangle has a base of 6 inches and the other is 10 inches. The total area of the two triangles is square inches.

AZ:

First, find the area of each triangle and then we add it up to get the total area of the two triangles The first triangle has a base of 6 inches and a height of 8 Area of a triangle = 1/2 * base * height what is the area of this triangle?

rxcklesskaisher:

48

AZ:

No, remember that it's 1/2 * base * height so what is 1/2 * 6 * 8 =

rxcklesskaisher:

24

AZ:

wait, my apologies, the height is 10 so what is 1/2 * 6 * 10 =

rxcklesskaisher:

30

AZ:

Good! Now let's find the area of the other triangle the base is 8 the height is 10 Area = 1/2 * base * height Area = ?

rxcklesskaisher:

30

AZ:

No, we have new numbers for the OTHER triangle base = 8 height = 10 area = 1/2 * 8 * 10 area = ??

rxcklesskaisher:

40

AZ:

Good So what is the area of both triangles? Just add them up 30 + 40 = ?

rxcklesskaisher:

70

AZ:

and so that's the total area of the two triangles